# EXFormer: A Multi-Scale Trend-Aware Transformer with Dynamic Variable Selection for Foreign Exchange Returns Prediction

### v1

@@ 3 Methodology
In this section, we present the EXFormer architecture in detail. The proposed EXFormer architecture is conceptually inspired by the trend-aware self-attention model [li2019enhancing, guo2021learning]. We have redesigned this foundation into a novel multi-scale trend-aware self-attention mechanism. The motivation for this multi-scale approach is to effectively capture the complex temporal patterns in the daily exchange rate of high-frequency fluctuations, medium-term cycles, and long-run drifts. Recognizing that the daily exchange rates are driven by numerous external factors, we have developed a dynamic variable selector that not only reweights variables dynamically but also imparts pre-hoc interpretability to the architecture. Meanwhile, to contend with the issue of information redundancy in financial time series forecasting, the model employs a squeeze-and-excitation (SE) block. The above is among the Encoder. For the Decoder, we are inspired by shu2025deformtime to introduce GRU and linear mapping. Figure 1 provides a schematic overview of the entire model.
@@ 3.3 Multi-scale convolution and squeeze-and-excitation
In line with prior studies on multi-scale convolution for financial and economic time series [liang2022airformer, pmlr-v162-lan22a, guo2021learning], we adopt 3 convolutional branches to capture short-, medium-, and long-horizon return dynamics. The kernel sizes are initialized as $\{3,5,7\}$ to reflect these scales, while their exact values are subsequently tuned automatically during hyperparameter optimization.
@@ 3.4 Multi-scale trend-aware self-attention
As suggested in related work on trend-aware attention mechanisms for time series [liang2022airformer, pmlr-v162-lan22a, guo2021learning], we design three parallel convolutional branches, each operating with a distinct kernel size to represent short-, medium-, and long-term dependencies. The convolution kernel sizes are initialized at $\{3,5,7\}$ following common practice, while the actual receptive fields are subject to automatic selection in the hyperparameter tuning process.
@@ 4.1 Data description
In addition, we assess volatility clustering in the daily returns using ARCH-LM tests. For all three currency pairs, the null hypothesis of no ARCH effects is rejected at conventional significance levels across multiple lags (i.e., LM-statistics exceeding 100 with $p<0.001$). The detailed test statistics are provided in the 15. This diagnostic addresses the concern that conditional heteroskedasticity may bias inference. In our empirical strategy, it is further mitigated by using heteroskedasticity and autocorrelation consistent (HAC) [newey1986simple] standard errors in forecast evaluation tests and by relying on risk-adjusted criteria in backtesting.
@@
We incorporate a broad set of external variables that capture information from various segments of the global financial and economic situation. Table 2 lists the set of covariates employed in this study, which are categorized into five groups: global foreign exchange, stock indices, fixed-income yields, commodity indices, and macroeconomic variables. These variables are chosen to provide a rich informational context, reflecting investor sentiment, monetary policy expectations, commodity price dynamics, and the overall health of the macroeconomy [candian2025imperfect, fang202430, mueller2017exchange, rime2010exchange]. The set includes other major exchange rates, key US stock indices like the S&P 500 and NASDAQ, a range of US Treasury yields from 3-months to 10-years, broad commodity indices, and macroeconomic indicators such as the economic policy uncertainty (EPU) index, CPI, and GDP growth. When forecasting one of the target exchange rates returns, the other two are included as covariates to capture the dynamic relationships within the foreign exchange market. For example, in the task of predicting EUR/USD returns, the daily returns of USD/JPY and GBP/USD are used as input features. To handle the different frequencies of the data, lower-frequency variables were forward-filled to align with the daily timescale, ensuring that no future information is leaked into the training process. All series spot exchange rates, equity indices, treasury yields, commodity indices, and macroeconomic indicators were retrieved from the Wind Database.
@@ 4.2 Experimental design and evaluation criteria
This study focus on 1 day-ahead forecasting of the exchange rate returns, the most challenging and practically relevant horizon for exchange rate market participants [darvas2025forecasting, engel2019uncovered, byrne2018sources], and assess performance across five look-back window lengths $T\in\{5,10,15,20,30\}$ trading days. The choice of these window lengths is motivated by the need to balance the amount of historical information used for prediction with the model’s ability to adapt to changing market conditions. The shorter windows capture recent trends and volatility, while the longer windows provide a broader context of historical behavior.
We compare EXFormer against popular deep learning architectures and a classical econometric model. The vanilla Transformer with identical embedding dimension and number of layers, to isolate the impact of our multi-scale trend-aware modifications. Temporal Convolutional Network (TCN) is a purely convolutional sequence model that achieves long receptive fields via dilated convolutions, representing an alternative multi-scale feature extractor. GRU and LSTM are two recurrent architectures widely used in financial time-series forecasting, capable of capturing sequential dependencies but lacking multi-scale or interpretability modules. Multilayer Perceptron (MLP) is a dense feed-forward network on lagged inputs, serving as a simple non-sequential benchmark. ARIMA is a univariate autoregressive integrated moving-average model, chosen for its frequent use in time-series analysis. Beyond these standard baselines, we also include: Informer, an attention-based architecture designed for long-sequence time-series forecasting with a probabilistic sparse self-attention mechanism that reduces computational complexity while preserving predictive power. LSTM-ARIMA [kashif2025lstm], a hybrid approach that integrates linear ARIMA components with nonlinear sequence modeling from LSTM, allowing the capture of both autoregressive structures and nonlinear dependencies; and the large language model (LLM) Chronos (chronos-t5-base) [ansari2024chronos], a transformer-based foundation model recently adapted for time-series forecasting, which serves as a cutting-edge benchmark reflecting the capabilities of pretrained LLMs applied to financial prediction.
@@
We assess predictive accuracy along three dimensions. First, following [darvas2025forecasting, rossi2013exchange], we compute the mean squared forecast error (MSFE) of each model and express it as a percentage relative to that of the random walk. For model $m$, the MSFE is defined as
@@
with values below $100$ indicating an improvement over the random walk (RW). To assess statistical significance in nested comparisons, we apply the Clark–West test [clark2007approximately], estimating HAC standard errors via newey1986simple estimator with Bartlett kernel. Clark–West’s correction accounts for the noise introduced by parameter estimation in the larger model.
@@
Here, $N$ is the number of forecasting instances, $\hat{y}_{i}$ is the predicted return, and $y_{i}$ is the true return. The values of DA above 0.5 indicate better-than-chance performance. To account for serial correlation in the indicator series when using multi-day horizons, we follow blaskowitz2014testing and estimate the dynamic regression using Newey–West HAC standard errors to obtain a robust $t$-statistic and $p$-value for the null that the model’s directional accuracy does not exceed that of the random walk.
@@ 6 Discussion
Despite a growing number of studies applying machine learning and neural networks to daily exchange rate returns forecasting, most have only compared their novel architectures against conventional baselines, such as simple neural networks or single factor models, and stopped short of benchmarking against the random walk or demonstrating statistical superiority over it [windsor2022improving, mao2024unveiling]. For example, prior works by fischer2018deep report backtest returns for their proposed networks but do not conduct formal statistical tests to confirm outperformance of the random walk. In contrast, EXFormer not only registers statistically significant gains in out-of-sample MSFE and directional accuracy, but also translates those gains into economically meaningful trading profits, outpacing buy and hold, moving average, and random walk strategies across multiple currency pairs. Our built-in interpretability mechanism yields additional economic insights. Whereas studies such as afshan2024fintech and singhal2019return highlight the predictive power of crude oil and gold markets, our global importance analysis also shows that broad commodity indices (e.g. S&P GSCI and Bloomberg Commodity) exert the largest influence on day-ahead exchange rate returns [beckmann2020relationship]. Moreover, the dynamic variable weights reveal that EUR/USD returns are particularly sensitive to the stock markets and U.S. Treasury yields, corroborating findings by ben2024shorting on the close coupling between the euro and global financial conditions.
@@
Furthermore, our sliding window analysis points to an empirical optimum at 15 trading days for 1 day-ahead exchange rate returns prediction. While some econometric studies (e.g. using AIC or cross-validation) recommend longer or varying look-back lengths[darvas2025forecasting], and others claim robustness across any reasonably sized window [byrne2018sources], our results suggest that information beyond three weeks begins to weaken the model’s ability to adapt to rapid market regime shifts. The finding has practical implications that careful tuning of the sliding window is as critical as neural network architecture when deploying deep learning for high-frequency financial forecasting. Moreover, foreign exchange is one of the most liquid asset classes globally, with deep order books and minimal transaction frictions at daily horizons [engel2023liquidity]. Therefore, liquidity constraints are unlikely to bias our predictive results or trading simulations, distinguishing this setting from equity or commodity markets where limited depth or slippage can materially affect forecast evaluation [erdogan2015recession].
@@
Our findings contribute to the longstanding debate on the validity of classical exchange rate theories. The consistent profitability of EXFormer in daily forecasting underscores the well-documented short-horizon failure of UIP [engel2019uncovered], often referred to as the forward premium puzzle, whereby exchange rates do not fully adjust to eliminate predictable excess returns. This result aligns with the extensive empirical evidence that UIP is systematically broke at daily horizons [candian2025imperfect, du2018deviations, gabaix2015international]. At the same time, the interpretability analysis highlights the role of global commodities, equity indices, and long-term interest rates in shaping currency dynamics. Such evidence corresponds with the portfolio balance perspective [ben2024shorting], which emphasizes the importance of cross-asset capital flows and shifts in investor risk preferences in determining exchange rates [verdelhan2018share, mueller2017exchange]. By jointly demonstrating UIP’s empirical breakdown and providing support for portfolio balance mechanisms, our study situates the predictive and economic value of EXFormer within established theoretical frameworks and offers a bridge between modern machine-learning methods and classical international finance theories.

### v2

@@ 3 Methodology
In this section, we present the EXFormer architecture in detail. The proposed EXFormer architecture is conceptually inspired by the trend-aware self-attention model [44, 33]. We have redesigned this foundation into a novel multi-scale trend-aware self-attention mechanism. The motivation for this multi-scale approach is to effectively capture the complex temporal patterns in the daily exchange rate of high-frequency fluctuations, medium-term cycles, and long-run drifts. Recognizing that the daily exchange rates are driven by numerous external factors, we have developed a dynamic variable selector that not only reweights variables dynamically but also imparts pre-hoc interpretability to the architecture. Meanwhile, to contend with the issue of information redundancy in financial time series forecasting, the model employs a squeeze-and-excitation (SE) block. The above is among the Encoder. For the Decoder, we are inspired by Shu and Lampos [66] to introduce GRU and linear mapping. Figure 1 provides a schematic overview of the entire model.
@@ 3.3 Multi-scale convolution and squeeze-and-excitation
In line with prior studies on multi-scale convolution for financial and economic time series [45, 43, 33], we adopt 3 convolutional branches to capture short-, medium-, and long-horizon return dynamics. The kernel sizes are initialized as $\{3,5,7\}$ to reflect these scales, while their exact values are subsequently tuned automatically during hyperparameter optimization.
@@ 3.4 Multi-scale trend-aware self-attention
As suggested in related work on trend-aware attention mechanisms for time series [45, 43, 33], we design three parallel convolutional branches, each operating with a distinct kernel size to represent short-, medium-, and long-term dependencies. The convolution kernel sizes are initialized at $\{3,5,7\}$ following common practice, while the actual receptive fields are subject to automatic selection in the hyperparameter tuning process.
@@ 4.1 Data description
In addition, we assess volatility clustering in the daily returns using ARCH-LM tests. For all three currency pairs, the null hypothesis of no ARCH effects is rejected at conventional significance levels across multiple lags (i.e., LM-statistics exceeding 100 with $p<0.001$). The detailed test statistics are provided in the 15. This diagnostic addresses the concern that conditional heteroskedasticity may bias inference. In our empirical strategy, it is further mitigated by using heteroskedasticity and autocorrelation consistent (HAC) [59] standard errors in forecast evaluation tests and by relying on risk-adjusted criteria in backtesting.
@@
We incorporate a broad set of external variables that capture information from various segments of the global financial and economic situation. Table 2 lists the set of covariates employed in this study, which are categorized into five groups: global foreign exchange, stock indices, fixed-income yields, commodity indices, and macroeconomic variables. These variables are chosen to provide a rich informational context, reflecting investor sentiment, monetary policy expectations, commodity price dynamics, and the overall health of the macroeconomy [12, 23, 56, 62]. The set includes other major exchange rates, key US stock indices like the S&P 500 and NASDAQ, a range of US Treasury yields from 3-months to 10-years, broad commodity indices, and macroeconomic indicators such as the economic policy uncertainty (EPU) index, CPI, and GDP growth. When forecasting one of the target exchange rates returns, the other two are included as covariates to capture the dynamic relationships within the foreign exchange market. For example, in the task of predicting EUR/USD returns, the daily returns of USD/JPY and GBP/USD are used as input features. To handle the different frequencies of the data, lower-frequency variables were forward-filled to align with the daily timescale, ensuring that no future information is leaked into the training process. All series spot exchange rates, equity indices, treasury yields, commodity indices, and macroeconomic indicators were retrieved from the Wind Database.
@@ 4.2 Experimental design and evaluation criteria
This study focus on 1 day-ahead forecasting of the exchange rate returns, the most challenging and practically relevant horizon for exchange rate market participants [18, 20, 10], and assess performance across five look-back window lengths $T\in\{5,10,15,20,30\}$ trading days. The choice of these window lengths is motivated by the need to balance the amount of historical information used for prediction with the model’s ability to adapt to changing market conditions. The shorter windows capture recent trends and volatility, while the longer windows provide a broader context of historical behavior.
We compare EXFormer against popular deep learning architectures and a classical econometric model. The vanilla Transformer with identical embedding dimension and number of layers, to isolate the impact of our multi-scale trend-aware modifications. Temporal Convolutional Network (TCN) is a purely convolutional sequence model that achieves long receptive fields via dilated convolutions, representing an alternative multi-scale feature extractor. GRU and LSTM are two recurrent architectures widely used in financial time-series forecasting, capable of capturing sequential dependencies but lacking multi-scale or interpretability modules. Multilayer Perceptron (MLP) is a dense feed-forward network on lagged inputs, serving as a simple non-sequential benchmark. ARIMA is a univariate autoregressive integrated moving-average model, chosen for its frequent use in time-series analysis. Beyond these standard baselines, we also include: Informer, an attention-based architecture designed for long-sequence time-series forecasting with a probabilistic sparse self-attention mechanism that reduces computational complexity while preserving predictive power. LSTM-ARIMA [41], a hybrid approach that integrates linear ARIMA components with nonlinear sequence modeling from LSTM, allowing the capture of both autoregressive structures and nonlinear dependencies; and the large language model (LLM) Chronos (chronos-t5-base) [3], a transformer-based foundation model recently adapted for time-series forecasting, which serves as a cutting-edge benchmark reflecting the capabilities of pretrained LLMs applied to financial prediction.
@@
We assess predictive accuracy along three dimensions. First, following [18, 63], we compute the mean squared forecast error (MSFE) of each model and express it as a percentage relative to that of the random walk. For model $m$, the MSFE is defined as
@@
with values below $100$ indicating an improvement over the random walk (RW). To assess statistical significance in nested comparisons, we apply the Clark–West test [17], estimating HAC standard errors via Newey and West [59] estimator with Bartlett kernel. Clark–West’s correction accounts for the noise introduced by parameter estimation in the larger model.
@@
Here, $N$ is the number of forecasting instances, $\hat{y}_{i}$ is the predicted return, and $y_{i}$ is the true return. The values of DA above 0.5 indicate better-than-chance performance. To account for serial correlation in the indicator series when using multi-day horizons, we follow Blaskowitz and Herwartz [7] and estimate the dynamic regression using Newey–West HAC standard errors to obtain a robust $t$-statistic and $p$-value for the null that the model’s directional accuracy does not exceed that of the random walk.
@@ 6 Discussion
Despite a growing number of studies applying machine learning and neural networks to daily exchange rate returns forecasting, most have only compared their novel architectures against conventional baselines, such as simple neural networks or single factor models, and stopped short of benchmarking against the random walk or demonstrating statistical superiority over it [78, 50]. For example, prior works by Fischer and Krauss [25] report backtest returns for their proposed networks but do not conduct formal statistical tests to confirm outperformance of the random walk. In contrast, EXFormer not only registers statistically significant gains in out-of-sample MSFE and directional accuracy, but also translates those gains into economically meaningful trading profits, outpacing buy and hold, moving average, and random walk strategies across multiple currency pairs. Our built-in interpretability mechanism yields additional economic insights. Whereas studies such as Afshan et al. [2] and Singhal et al. [67] highlight the predictive power of crude oil and gold markets, our global importance analysis also shows that broad commodity indices (e.g. S&P GSCI and Bloomberg Commodity) exert the largest influence on day-ahead exchange rate returns [5]. Moreover, the dynamic variable weights reveal that EUR/USD returns are particularly sensitive to the stock markets and U.S. Treasury yields, corroborating findings by Ben Zeev and Nathan [6] on the close coupling between the euro and global financial conditions.
@@
Furthermore, our sliding window analysis points to an empirical optimum at 15 trading days for 1 day-ahead exchange rate returns prediction. While some econometric studies (e.g. using AIC or cross-validation) recommend longer or varying look-back lengths[18], and others claim robustness across any reasonably sized window [10], our results suggest that information beyond three weeks begins to weaken the model’s ability to adapt to rapid market regime shifts. The finding has practical implications that careful tuning of the sliding window is as critical as neural network architecture when deploying deep learning for high-frequency financial forecasting. Moreover, foreign exchange is one of the most liquid asset classes globally, with deep order books and minimal transaction frictions at daily horizons [21]. Therefore, liquidity constraints are unlikely to bias our predictive results or trading simulations, distinguishing this setting from equity or commodity markets where limited depth or slippage can materially affect forecast evaluation [22].
@@
Our findings contribute to the longstanding debate on the validity of classical exchange rate theories. The consistent profitability of EXFormer in daily forecasting underscores the well-documented short-horizon failure of UIP [20], often referred to as the forward premium puzzle, whereby exchange rates do not fully adjust to eliminate predictable excess returns. This result aligns with the extensive empirical evidence that UIP is systematically broke at daily horizons [12, 19, 30]. At the same time, the interpretability analysis highlights the role of global commodities, equity indices, and long-term interest rates in shaping currency dynamics. Such evidence corresponds with the portfolio balance perspective [6], which emphasizes the importance of cross-asset capital flows and shifts in investor risk preferences in determining exchange rates [74, 56]. By jointly demonstrating UIP’s empirical breakdown and providing support for portfolio balance mechanisms, our study situates the predictive and economic value of EXFormer within established theoretical frameworks and offers a bridge between modern machine-learning methods and classical international finance theories.